A 5-year loan in the amount of $48,000 is to be repaid in equal annual payments. What is the remaining principal balance after the third payment if the interest rate is 5 percent, compounded annually?
Answer #1
Rate Annual Interest rate 5%
Nper Number of Years 5
Pv Loan amount $48,000
PMT Equal Annual Payment $11,086.79 (Using PMT function of excel with Rate=5%,Nper=5,Pv=-48000)
Principal Balance after 3rd payment=Present Value of future payments
Rate Annual Interest rate 5%
Nper Number of Years left in future=5-3 2
Pmt Annual payment $11,086.79
PV Principal Balance after 3rd payment $20,614.89 (Using PV function of excel with Rate=5%,Nper=2,Pmt=-11086.79)
SCHEDULE OF PAYMENTS A B C=A*5% D=B-C E=A-D
Year Beginning Loan Balance Loan Repayment Interest Principal Ending Loan Balance
0 $48,000
1 $48,000 $11,086.79 $2,400 $8,686.79 $39,313.21
2 $39,313 $11,086.79 $1,966 $9,121.13 $30,192.08
3 $30,192 $11,086.79 $1,510 $9,577.19 $20,614.89
4 $20,615 $11,086.79 $1,031 $10,056.05 $10,558.85
5 $10,559 $11,086.79 $528 $10,558.85 $0.00

D6 x fc =PMT(D3,D4,-05) D E F G H I J 5% 2 3 Rate 4 Nper 5 PV 6 PMT Annual Interest rate Number of Years Loan amount Equal An012 : X fic =PV(09,010,-011) D B E F G H I 5% Annual Interest rate Number of Years Loan amount Equal Annual Payment $48,000 $